OpenStudy (anonymous):

Abby is registering at a Web site. She must select a password containing 6 numerals to be able to use the site. How many passwords are allowed if no digit may be used more than once?

OpenStudy (anonymous):

10P6 = 10!/(10-6)! = 10!/4! = 10*9*8*7*6*5

OpenStudy (anonymous):

10*9*8*7*6*5 = 151,200

OpenStudy (anonymous):

so 151200

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh ok thanks